OpenStudy (anonymous):

A falling stone takes 0.2 seconds to fall past a window which is 1 m high. From how far above the top of the window was the stone dropped?

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

shall i show the working?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

s= ut +1/2at^2 s- 1 = 1/2 * 9.81 * (0.2)^2 s = 1.196

OpenStudy (anonymous):

s=ut+1/2 at^2 x=0+1/2 * 9.81 * (0.2)^2 x=0.196 m

OpenStudy (anonymous):

Is the answer 0.196m

OpenStudy (anonymous):

but the answer is 0.82

OpenStudy (anonymous):

oh got it..

OpenStudy (anonymous):

Actually, s= ut + 1/2 at^2 1 = u*0.2 + 1/2 (9.81) *(0.2^2) u=4.02 m/s

OpenStudy (anonymous):

a falling stone right

OpenStudy (anonymous):

Now, v^2=u^2+2as (4.02)^2=0^2+2*(9.81)*x x=0.82m

OpenStudy (anonymous):

Yes, the falling stone takes 0.2 seconds to reach the ground from the window

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):